I need help on #18 please.

3.15 uF 4.4 F 7.88 UF 누 3.56 UF 10.8 V 는

Find the charge on 7.88 uF capacitor on the right. Answer in units of uc.
Answer #1

We have \small C_{1}=4.4\mu f ,\small C_{2}=3.15\mu f\, ,\, C_{3}=3.56\mu f,\, C_{4}=7.88\mu f

First we will simply the circuit and find equivalent capacitance of \small C_{2}=3.15\mu f\, and\, C_{3}=3.56\mu f\,

since they are parallel to each other \small C^{'}_{eq}= C_{2} + C_{3}=3.15+3.56=6.71\mu f

Now all three capacitors \small C^{'}_{eq},C_{1}, C_{4} are in series,hence have same charge in all of them.

NOW we calculate equivalent capacitance of complete circuit.

\small 1/C_{eq}=1/C^{'}_{eq}+1/C_{1}+1/ C_{4} = (1/6.71)+(1/4.4)+(1/7.88)

hence \small C_{eq}=1.98\mu f

since all capacitor having same charge,using Q=CV=1.98x10.8=\small 21.4\mu C

using same formula for \small C=7.88\mu f , V=Q/C= 21.4/7.88= 2.72V

hence V = 2.72V